Solve the inequality 2|x-10|less than sign 3, and express the solution as an interval. Thank you!
Answer by venugopalramana(3286)

2|X-10|<3
|X-10|<3/2=1.5
CASE 1...X-10 IS +VE
|X-10|=X-10<1.5
X<11.5
CASE2....X-10 IS -VE
|X-10|=-X+10<1.5
-X<-8.5
X>8.5
HENCE 8.5 OR X IS (8.5,11.5)